A cylinder contains 400g of oxygen and 600g of helium at a total pressure of 7.0 atm. Calculate partial pressures of the gases.

1
Expert's answer
2022-01-21T08:52:04-0500

Mole ratio : O:He

=2:3

Partial pressure of oxygen =( 2/ 5 )× 7= 2.8 ATM

Partial pressure of helium=( 3/5) × 7= 4.2 atm